OpenStudy (anonymous):

Caleb divided all of his CDs onto 3 shelves. Each shelf holds twelve CDs. Caleb wants to know the total number of CDs he has. Which of the following is a verbal model for this problem? number of shelves divided by total CDs equals number of CDs on each shelf number of CDs on each shelf times total CDs equals number of shelves number of shelves divided by number of CDs on each shelf equals total CDs number of shelves times number of CDs on each shelf equals total CDs

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

It would be the last option.

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

What is the first step in solving the equation eight d minus six equals two for d? Add 6 Subtract 6 Divide by 8 Multiply by 8

OpenStudy (anonymous):

can someone help

OpenStudy (anonymous):

8d-6=2 First we add six to both sides 8d=8 Then we get d by itself by dividing both sides by 8 d=1

OpenStudy (anonymous):

Which equation best represents the following scenario? Heidi answered 4 more questions correctly on the quiz than Sean. Heidi answered 8 questions correctly. four s equals eight s plus four equals eight four minus s equals eight eight divided by s equals four

OpenStudy (anonymous):

s+4=8 s=4

OpenStudy (anonymous):

Which inequality is represented by this graph? A number line shows the numbers zero to ten. There is an open dot at nine, and an arrow extending from that open dot toward the left. three x is less than twenty-seven x divided by three is less than or equal to three four x is greater than or equal to thirty-six x divided by three is less than nine

OpenStudy (anonymous):

@MChromer can you help me

OpenStudy (anonymous):

It is the last option because the dot is at nine and pointing to numbers less than nine.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Which of the following equations does not require two steps to solve? negative seven b minus thirteen equals one sixteen minus r equals eight negative thirty-two plus three k equals negative twenty-six w minus twenty-two equals negative fifteen

OpenStudy (anonymous):

It is the last one, it only takes one step.

OpenStudy (anonymous):

Which property would you use to solve the equation zero point seven p equals zero point two one for p? Addition Property of Equality Multiplication Property of Equality Subtraction Property of Equality Division Property of Equality

OpenStudy (anonymous):

I don't knowproperties very well, but I'm guessing Division Property of Equality

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

What is the first step in solving the equation seven d plus one equals eight for d?

OpenStudy (anonymous):

d+1=8 subtract one d=7